What is the molarity of a solution when 157g of CaSO4 are dissolved in enough water to yield a volume of 724ml
1
Expert's answer
2021-08-11T05:21:52-0400

M(CaSO4) = 136.14 g/mol

"n=\frac{m}{M} \\

= \frac{157}{136.14}=1.153 \;mol"

Proportion:

1.153 mol – 724 mL

x mol – 1000 mL

"x = \frac{1.153 \times 1000}{724}= 1.59 \;mol/L ≈1.6 \;M"

Answer: 1.6 M